2 LINEAR EQUATIONS IN 2 UNKNOWNS or SIMULTANEOUS LINEAR EQUATIONS IN 2 UKNOWNS
The sum of the digits of a two digit number is 14. When the digit are reversed, the resulting number is 23 less than two times the original number. What is the number?

Answer by rothauserc(4718)

1) x +y = 14
:
2) 10y +x = 2(10x +y) -23
:
simplify equation 2
:
10y +x = 20x +2y -23
:
3) 19x -8y = 23
:
solve equation 1 for x and substitute in equation 3
:
19(14-y) -8y = 23
:
266 -19y -8y = 23
:
-27y = -243
:
y = 9
:
x = 14 -9 = 5
:
************************************
x = 5 and y = 9 and the number is 59
:
check the answer
:
95 = 2(59) -23
:
95 = 95
*************************************
: